How is the Cash Conversion Cycle of the fashion retailer Nordstrom be different from that of Walmart?
Answer #1

The Cash conversion cycle is calculated using days sales outstanding , days inventory and days payable.

CCC = DSO + DI - DP

Nordstorm is a fashion retailer who are expected to have a longer inventory period, which eventually leads to longer days payable, because if we have inventory for long , we will receive money late , so we will pay late.

DSO of Nordstorm for 0ct 2019 = 5.81

DI = 87.08

DP = 83.62

Therefore,

CCC = 5.81 + 87.08 - 83.62 = 9.21

Walmart is a retail corporation that operates a chain of hypermarkets , discount department stores , and grocery stores. The inventory in these companies stays for a lesser time period than fashion companies , because the product are cheap and required on daily basis. This helps them to get money faster than the fashion company and pay creditors at a faster rate.

DSO of walmart for October 2019 = 4

DI = 45.52

DP = 47.34

CCC = 4 + 45.52 - 47.34 = 2.18

We can see from the figures, that the cash conversion cycle of walmart is less than Nordstorm. The main difference which we can see is in the days inventory and days payable, as explaned above. This shows how a fashion industry and a grocery and daily use products company work.